Madrano's Wholesale Fruit Company located in​ McAllen, Texas is considering the purchase of a new fleet of tractors to be used in the delivery of fruits and vegetables grown in the Rio Grande Valley of Texas. If it goes through with the​ purchase, it will spend $300,000. on eight rigs. The new trucks will be kept for 5 years, during which time they will be depreciated toward a $43,000 salvage value using​ straight-line depreciation. The rigs are expected to have a market value in 5 years equal to their salvage value. The new tractors will be used to replace the​ company's older fleet of eight trucks which are fully depreciated but can be sold for an estimated $18,000 (because the tractors have a current book value of​ zero, the selling price is fully taxable at the​ firm's 25 percent tax​ rate). The existing tractor fleet is expected to be useable for 5 more years after which time they will have no salvage value. The existing fleet of tractors uses $210,000 per year in diesel​ fuel, whereas the​ new, more efficient fleet will use only $110,000 in​ addition, the new fleet will be covered under​ warranty, so the maintenance costs per year are expected to be only $13,000 compared to $32,000 for the existing fleet.

a.  What are the differential operating cash flow savings per year during years 1 through 5 for the new​ fleet?

b.  What is the initial cash outlay required to replace the existing fleet with the newer​ tractors?

c.  What does the timeline for the replacement project cash flows for years 0 through 5 look like?

d.  If Madrano requires a discount rate of 11% for new​ investments, should the fleet be​ replaced?
Answer #1

Answer Page No ① Opertaling cash blows www Fuel net of tax = 210.000 X61-25%) firms for reade a = 25% . 210,000 X 0.75 = $154Page No @ Depreciation - 300,000 - 43000 5 years 257000 = $51400 ... Tar saving on depreciation - 51400 X 25% = $ 12850 NewsPage No ③ Feitical cases ootan erequired Initial cash outlay srequired = cool g news feet – Selvoge value of old net of taxePage No 4 = 102100 X PVA FC 11%, 5yrs) + 42000 4 PUF (P8%95995) - 286500 = (109100 x Chiku 299(1280 x cu mest) - 286500 (102